# A low-rank solver for conforming multipatch Isogeometric Analysis

### v3

@@ 3.3 The TPCG method with block-wise truncation
The simplest truncation operator considered here is relative truncation and it is denoted with $\mathsf{{T}^{rel}}$. Given a Tucker vector $\boldsymbol{v}$ and a relative tolerance $\epsilon>0$, $\boldsymbol{\tilde{v}}=\mathsf{{T}^{rel}}(\boldsymbol{v},\eta)$ is a Tucker vector with smaller or equal multilinear rank (in each direction) such that
$\|\boldsymbol{\tilde{v}}-\boldsymbol{v}\|_{2}\leq\epsilon\|\boldsymbol{v}\|_{2}.$ (3.8)
We refer to [21, Section 4.1.1] for more details.
We also consider a dynamic truncation operator $\mathsf{{T}^{dt}}$, which is used only for the approximate solution $\boldsymbol{u}_{k+1}$. Here a truncated approximation $\boldsymbol{\tilde{u}}_{k+1}$ of $\boldsymbol{u}_{k+1}$ is initially computed using block-wise relative truncation with a prescribed initial tolerance $\epsilon$. Then the algorithm assesses how much the exact solution update $\Delta\boldsymbol{u}_{k}=\boldsymbol{u}_{k+1}-\boldsymbol{u}_{k}$ differs from the truncated one $\Delta\boldsymbol{\tilde{u}}_{k}=\boldsymbol{\tilde{u}}_{k+1}-\boldsymbol{u}_{k}$ by checking if the condition
$\left|\frac{\Delta\boldsymbol{\tilde{u}}_{k}\cdot\Delta\boldsymbol{u}_{k}}{\|\Delta\boldsymbol{u}_{k}\|^{2}_{2}}-1\right|\leq\delta$
is satisfied, where $\delta>0$ is a prescribed tolerance. If the condition is satisfied, then $\boldsymbol{\tilde{u}}_{k+1}$ is accepted as the next iterate. Otherwise, a new tolerance for the relative truncation is selected, equal to $\max\left\{\alpha\epsilon,\epsilon_{\min}\right\}$, where $10<\alpha<$ 1 and $\epsilon_{\min}>0$ is a fixed minimum tolerance. The process is repeated until an acceptable truncated solution is found (or until the tolerance reaches $\epsilon_{\min}$). The final relative tolerance is returned and it is used as initial tolerance for $\mathsf{{T}^{dt}}$ at the next TPCG iteration. A starting relative tolerance $\epsilon_{0}$ is fixed at the beginning of the TPCG algorithm. We refer to [21, Section 4.1.2] for more details.
Except for $\boldsymbol{\tilde{u}}_{k+1}$, all other vectors are compressed using block-wise relative truncation with tolerance
@@
where $\boldsymbol{r}_{0}$ and $\boldsymbol{r}_{k}$ are the initial and $k-$th residual vector, respectively, $10<\beta<$ 1 and tol is the TPCG tolerance. This choice of the tolerance promotes low rank, see also [28, 23]. Furthermore, we introduce intermediate truncation steps during the computation of matrix-vector products with $\mathbf{A}$, motivated by the significant rank increment yielded by this operation. We remark, however, that intermediate truncation can lead to cancellation errors and stagnation, see e.g. the discussion in [29, Section 3.6.3]. Therefore, a strict relative tolerance $\gamma$ should be chosen to safely perform this step.
@@
If the magnitude of the solution varies significantly in different subdomains, it is a good idea to perform truncations on the vector blocks such that the absolute error (instead of the relative error) is equidistributed. This help in balancing the error among different subdomains. Therefore, given a vector block $\boldsymbol{y}^{(i,k)}$, belonging to a vector $\boldsymbol{y}$, Its truncation $\tilde{\boldsymbol{y}}^{(i,k)}$ could be chosen by imposing a condition of the form
$\|\boldsymbol{\tilde{y}}^{(i,k)}-\boldsymbol{y}^{(i,k)}\|_{2}\leq\frac{\eta}{\mathcal{N}_{sub}}\|\boldsymbol{y}\|_{2},$
@@ 3.4 Subdomains as unions of patches
In the setting described in Section 2.2, we can choose the subdomains as unions of pairs of patches sharing a face. More precisely, for each face shared by two patches, we introduce a subdomain defined as the union of those patches. Note that in this case $\mathcal{N}_{sub}$ is equal to the number of faces belonging to the interface $\Gamma$.
@@
$V_{sub}^{(i)}=\left\{\hat{v}_{h}\circ\mathcal{G}_{i}^{-1}\middle|\;\hat{v}_{h}\in\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}\right\}.$
@@
In the case of arbitrary boundary conditions, merging two local spline spaces does not necessarily results in a space which is the pushforward of a tensor product spline space. In this case, one can define $V_{sub}^{(i)}$ by substituting any Neumann boundary conditions with homogenous Dirichlet. Then an additional subdomain should be introduced containing the Neumann degrees of freedom.
@@
We now discuss in detail how the approximations (3.5) and (3.6) can be performed using low-rank techniques. We first consider the case of a matrix block $\mathbf{A}^{(i,j)}$ for $i\neq j$, $i,j\in\left\{1,\ldots,\mathcal{N}_{sub}\right\}$. Of course if $\Theta^{(i)}\cap\Theta^{(j)}$ has an empty interior, then $\mathbf{A}^{(i,j)}$ is a null matrix. On the other hand, if $\Theta^{(i)}\cap\Theta^{(j)}$ has a non-empty interior, then there is an index $m=m(i,j)\in\left\{1,\ldots,\mathcal{N}_{ptc}\right\}$ such that $\Theta^{(i)}\cap\Theta^{(j)}=\Omega^{(m)}$. Let $v_{h}\in V^{(i)}_{sub}$ and $w_{h}\in V^{(j)}_{sub}$, and let $\boldsymbol{v}$ and $\boldsymbol{w}$ be their vector representations. Then for $k,\ell=1,2,3$, it holds
@@
$\begin{split}\boldsymbol{v}^{T}\mathbf{A}^{(i,j,k,\ell)}\boldsymbol{w}&=\int_{\Omega^{(m)}}2\mu\;\varepsilon(\underline{e}_{k}v_{h}):\varepsilon(\underline{e}_{\ell}w_{h})+\lambda\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\left(\nabla\cdot\underline{e}_{\ell}w_{h}\right)\mathrm{d}{\underline{x}}=\\ &=\int_{[0,1]^{3}}\left(\nabla\hat{v}_{ptc}^{(m)}\right)^{T}C^{(m,k,\ell)}\nabla\hat{w}_{ptc}^{(m)}\mathrm{d}{\underline{\xi}},\end{split}$ (3.10)
where $\hat{v}_{ptc}^{(m)}=v_{h}\circ\mathcal{F}_{m},\hat{w}_{ptc}^{(m)}=w_{h}\circ\mathcal{F}_{m}$, and
$C^{(m,k,\ell)}=|\det\left(J_{\mathcal{F}_{m}}\right)|J_{\mathcal{F}_{m}}^{-1}\left[\mu\left(\delta_{k,\ell}I_{3}+\underline{e}_{\ell}\underline{e}_{k}^{T}\right)+\lambda\underline{e}_{k}\underline{e}_{\ell}^{T}\right]J_{\mathcal{F}_{m}}^{-T}.$
@@
Here $\delta_{k,\ell}$ is the Kronecker delta. Note that $\hat{v}_{ptc}^{(m)},\hat{w}_{ptc}^{(m)}\in\widehat{\boldsymbol{\mathcal{S}}}_{ptc,0}^{(m)}$. Moreover, if $\text{supp}\left(v_{h}\right)\cap\Omega^{(m)}$ (respectively $\text{supp}\left(w_{h}\right)\cap\Omega^{(m)}$) has an empty interior then $\hat{v}_{ptc}^{(m)}$ (respectively $\hat{w}_{ptc}^{(m)}$) is the null function and $\boldsymbol{v}^{T}\mathbf{A}^{(i,j,k,\ell)}\boldsymbol{w}=0$.
We consider a low-rank approximation of the entries of $C^{(m,k,\ell)}$:
$C^{(m,k,\ell)}_{\alpha,\beta}\left(\xi_{1},\xi_{2},\xi_{3}\right)\approx\sum_{r_{1},r_{2},r_{3}}\mathfrak{c}_{r_{1},r_{2},r_{3}}^{(m,k,\ell,\alpha,\beta)}c^{(m,k,\ell,\alpha,\beta)}_{(1,r_{1})}\left(\xi_{1}\right)c^{(m,k,\ell,\alpha,\beta)}_{(2,r_{2})}\left(\xi_{2}\right)c^{(m,k,\ell,\alpha,\beta)}_{(3,r_{3})}\left(\xi_{3}\right).$ (3.11)
Plugging (3.11) into (3.10), we obtain
$\boldsymbol{v}^{T}\mathbf{A}^{(i,j,k,\ell)}\boldsymbol{w}\approx\sum_{r_{1},r_{2},r_{3}}\sum_{\alpha,\beta=1}^{3}\mathfrak{c}_{r_{1},r_{2},r_{3}}^{(m,k,\ell,\alpha,\beta)}\int_{[0,1]^{3}}c^{(m,k,\ell,\alpha,\beta)}_{(1,r_{1})}\;c^{(m,k,\ell,\alpha,\beta)}_{(2,r_{2})}\;c^{(m,k,\ell,\alpha,\beta)}_{(3,r_{3})}\;\partial_{\alpha}\hat{v}_{ptc}^{(m)}\;\partial_{\beta}\hat{w}_{ptc}^{(m)}\;d\underline{\xi}.$
@@
Therefore, thanks to the tensor structure of the spline space $\widehat{\boldsymbol{\mathcal{S}}}_{ptc,0}^{(m)}$ the matrix block $\mathbf{A}^{(i,j,k,\ell)}$ can be approximated as
$\mathbf{A}^{(i,j,k,\ell)}\approx\sum_{r_{1},r_{2},r_{3}}\sum_{\alpha,\beta=1}^{3}\mathfrak{c}_{r_{1},r_{2},r_{3}}^{(m,k,\ell,\alpha,\beta)}\widetilde{A}^{(i,j,k,\ell,\alpha,\beta)}_{(3,r_{3})}\otimes\widetilde{A}^{(i,j,k,\ell,\alpha,\beta)}_{(2,r_{2})}\otimes\widetilde{A}^{(i,j,k,\ell,\alpha,\beta)}_{(1,r_{1})},$ (3.12)
@@
There is of course more than one approach to compute the low-rank approximation (3.11). A suitable possibility, proposed in [20], is to interpolate the left-hand side function using a coarse spline space and then performing a High Order Singular Value Decomposition (HOSVD) [7]. In this work, however, we follow [21] and employ the strategy described in [9, 8], which is based on interpolation with Chebyshev polynomials.
@@
where $\hat{f}_{k}^{(i)}=\det\left(J_{\mathcal{G}_{i}})\right)\left(f_{k}\circ\mathcal{G}_{i}^{-1}\right)$ and $\hat{v}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}^{-1}\in\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}$. We consider a low-rank approximation
$\hat{f}_{k}^{(i)}\left(\xi_{1},\xi_{2},\xi_{3}\right)\approx\sum_{r_{1},r_{2},r_{3}}\mathfrak{f}^{(i,k)}_{r_{1},r_{2},r_{3}}f^{(i,k)}_{(1,r_{1})}\left(\xi_{1}\right)f^{(i,k)}_{(2,r_{2})}\left(\xi_{2}\right)f^{(i,k)}_{(3,r_{3})}\left(\xi_{3}\right)$ (3.14)
Plugging (3.14) into (3.13) we obtain
[…]
@@
We now discuss the technical details of the low-rank approximation (3.14) considered in this work. Assuming that $\mathcal{G}_{i}$ is given by (3.9), we first use the Chebfun3F algorithm [8] to independently compute low-rank approximations of $\hat{f}_{k}^{(i)}=\det\left(J_{\mathcal{G}_{i}}\right)\left(f_{k}\circ\mathcal{G}_{i}^{-1}\right)$ for $0\leq\xi_{3}\leq\frac{1}{2}$ and for $\frac{1}{2}<\xi_{3}\leq 1$. The reason for this separation is that $J_{\mathcal{G}_{i}}$ is possibly discontinuous at $\xi_{3}=\frac{1}{2}$ (corresponding to the interface between the two patches in the physical domain) and this could hinder the effectiveness of the considered low-rank Chebyshev approximation. Note that also $f_{k}$ could be discontinuous. We extend the two approximations of $\hat{f}_{k}^{(i)}$, denoted with $\tilde{f}_{k,1}^{(i)}$ and $\tilde{f}_{k,2}^{(i)}$, by zero outside of their respective domain. Then we consider
$\hat{f}_{k}^{(i)}\approx\tilde{f}_{k}^{(i)}=\tilde{f}_{k,1}^{(i)}+\tilde{f}_{k,2}^{(i)}.$
Note that the multilinear rank of $\tilde{f}_{k}^{(i)}$ is given by the sum of the multilinear ranks of $\tilde{f}_{k,1}^{(i)}$ and $\tilde{f}_{k,2}^{(i)}$. As a final step, we reduce the rank of the approximation using a purely algebraic approach based on the HOSVD.
@@
$\begin{split}\boldsymbol{v}^{T}\mathbf{A}^{(i,i,k,\ell)}\boldsymbol{w}&=\int_{\Theta^{(i)}}2\mu\;\varepsilon(\underline{e}_{k}v_{h}):\varepsilon(\underline{e}_{\ell}w_{h})+\lambda\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\left(\nabla\cdot\underline{e}_{\ell}w_{h}\right)\mathrm{d}{\underline{x}}=\\ &=\int_{[0,1]^{3}}\left(\nabla\hat{v}_{sub}^{(i)}\right)^{T}Q^{(i,k,\ell)}\nabla\hat{w}_{sub}^{(i)}\mathrm{d}{\underline{\xi}},\end{split}$ (3.15)
where $\hat{v}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}$, $\hat{w}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}\in\widehat{\boldsymbol{\mathcal{S}}}^{(i)}_{sub,0}$, and

### v4

@@ 3.3 The TPCG method with block-wise truncation
The simplest truncation operator considered here is relative truncation and it is denoted with $\mathsf{{T}^{rel}}$. Given a Tucker vector $\boldsymbol{v}$ and a relative tolerance $\epsilon>0$, $\boldsymbol{\widetilde{v}}=\mathsf{{T}^{rel}}(\boldsymbol{v},\eta)$ is a Tucker vector with smaller or equal multilinear rank (in each direction) such that
$\|\boldsymbol{\widetilde{v}}-\boldsymbol{v}\|_{2}\leq\epsilon\|\boldsymbol{v}\|_{2}.$ (3.10)
We refer to [34, Section 4.1.1] for more details.
We also consider a dynamic truncation operator $\mathsf{{T}^{dt}}$, which is used only for the approximate solution $\boldsymbol{u}_{k+1}$. Here a truncated approximation $\boldsymbol{\widetilde{u}}_{k+1}$ of $\boldsymbol{u}_{k+1}$ is initially computed using block-wise relative truncation with a prescribed initial tolerance $\epsilon$. Then the algorithm assesses how much the exact solution update $\Delta\boldsymbol{u}_{k}=\boldsymbol{u}_{k+1}-\boldsymbol{u}_{k}$ differs from the truncated one $\Delta\boldsymbol{\widetilde{u}}_{k}=\boldsymbol{\widetilde{u}}_{k+1}-\boldsymbol{u}_{k}$ by checking if the condition
$\left|\frac{\Delta\boldsymbol{\widetilde{u}}_{k}\cdot\Delta\boldsymbol{u}_{k}}{\|\Delta\boldsymbol{u}_{k}\|^{2}_{2}}-1\right|\leq\delta$
is satisfied, where $\delta>0$ is a prescribed tolerance. If the condition is satisfied, then $\boldsymbol{\widetilde{u}}_{k+1}$ is accepted as the next iterate. Otherwise, a new tolerance for the relative truncation is selected, equal to $\max\left\{\alpha\epsilon,\epsilon_{\min}\right\}$, where $10<\alpha<$ 1 and $\epsilon_{\min}>0$ is a fixed minimum tolerance. The process is repeated until an acceptable truncated solution is found (or until the tolerance reaches $\epsilon_{\min}$). The final relative tolerance is returned and it is used as initial tolerance for $\mathsf{{T}^{dt}}$ at the next TPCG iteration. A starting relative tolerance $\epsilon_{0}$ is fixed at the beginning of the TPCG algorithm. We refer to [34, Section 4.1.2] for more details.
Except for $\boldsymbol{\widetilde{u}}_{k+1}$, all other vectors are compressed using block-wise relative truncation with tolerance
@@
where $\boldsymbol{r}_{0}$ and $\boldsymbol{r}_{k}$ are the initial and $k-$th residual vector, respectively, $10<\beta<$ 1 and tol is the TPCG tolerance. This choice of the tolerance promotes low rank, see also [43, 38]. Furthermore, we introduce intermediate truncation steps during the computation of matrix-vector products with $\mathbf{A}$, motivated by the significant rank increment yielded by this operation. We remark, however, that intermediate truncation can lead to cancellation errors and stagnation, see e.g. the discussion in [45, Section 3.6.3]. Therefore, a strict relative tolerance $\gamma$ should be chosen to safely perform this step.
@@
If the magnitude of the solution varies significantly in different subdomains, it is a good idea to perform truncations on the vector blocks such that the absolute error (instead of the relative error) is equidistributed. This help in balancing the error among different subdomains. Therefore, given a vector block $\boldsymbol{y}^{(i,k)}$, belonging to a vector $\boldsymbol{y}$, Its truncation $\widetilde{\boldsymbol{y}}^{(i,k)}$ could be chosen by imposing a condition of the form
$\|\boldsymbol{\widetilde{y}}^{(i,k)}-\boldsymbol{y}^{(i,k)}\|_{2}\leq\frac{\eta}{\mathcal{N}_{sub}}\|\boldsymbol{y}\|_{2},$
@@ 3.4 Subdomains as unions of patches
In the setting described in Section 2.2, we can choose the subdomains as unions of patches. More precisely, the subdomains can be defined as follows. For every corner that is shared by 8 patches, we introduce a subdomain defined as the union of those 8 patches. Then, for every edge that it is shared by 4 patches and that is not contained in the interior of the previous subdomains, we introduce a new subdomain defined as the union of those 4 patches. Finally, for every face that is shared by 2 patches and that is not contained in the interior of the previous subdomains, we introduce a new subdomain defined as the union of those 2 patches. Note that this procedure guarantees that (3.1) is satisfied.
In the reminder of this section we give details on the function subspaces and on the construction of the Galerkin matrices for the case of subdomains defined as the union of pairs of patches. The treatment of other subdomains represents a straightforward extension of this case. Indeed, a 4-patches subdomain can be treated as the union of a pair of 2-patches subdomains. Similarly, an 8-patches subdomain can be treated as the union of a pair of 4-patches subdomains.
@@
$V_{sub}^{(i)}=\left\{\widehat{v}_{h}\circ\mathcal{G}_{i}^{-1}\middle|\;\widehat{v}_{h}\in\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}\right\}.$
@@
In the case of arbitrary boundary conditions, merging two local spline spaces does not necessarily results in a space which is the pushforward of a tensor product spline space (see some examples for the L-shaped domain in Figure 1). In this case, one can define $V_{sub}^{(i)}$ by substituting any Neumann boundary conditions with homogeneous Dirichlet. Then an additional subdomain should be introduced containing the Neumann degrees of freedom.
Figure 1: Examples of boundary conditions that are not allowed for the L-shaped domain. A remedy is proposed in Remark 2
@@
We now discuss in detail how the approximations (3.6) and (3.7) can be performed using low-rank techniques. We first consider the case of a matrix block $\mathbf{A}^{(i,j)}$ for $i\neq j$, $i,j\in\left\{1,\ldots,\mathcal{N}_{sub}\right\}$. Of course if $\Theta^{(i)}\cap\Theta^{(j)}$ has an empty interior, then $\mathbf{A}^{(i,j)}$ is a null matrix. On the other hand, if $\Theta^{(i)}\cap\Theta^{(j)}$ has a non-empty interior, then there is an index $m\in\left\{1,\ldots,\mathcal{N}_{ptc}\right\}$ such that $\Theta^{(i)}\cap\Theta^{(j)}=\Omega^{(m)}$. Let $v_{h}\in V^{(i)}_{sub}$ and $w_{h}\in V^{(j)}_{sub}$, and let $\boldsymbol{v}$ and $\boldsymbol{w}$ be their vector representations. Then for $k,\ell=1,2,3$, it holds
@@
$\begin{split}\boldsymbol{v}^{T}\mathbf{A}^{(i,j,k,\ell)}\boldsymbol{w}&=\int_{\Omega^{(m)}}2\mu\;\varepsilon(\underline{e}_{k}v_{h}):\varepsilon(\underline{e}_{\ell}w_{h})+\lambda\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\left(\nabla\cdot\underline{e}_{\ell}w_{h}\right)\mathrm{d}{\underline{x}}=\\ &=\int_{[0,1]^{3}}\left(\nabla\widehat{v}_{ptc}^{(m)}\right)^{T}C^{(m,k,\ell)}\nabla\widehat{w}_{ptc}^{(m)}\mathrm{d}{\underline{\xi}},\end{split}$ (3.12)
where $\widehat{v}_{ptc}^{(m)}=v_{h}\circ\mathcal{F}_{m},\widehat{w}_{ptc}^{(m)}=w_{h}\circ\mathcal{F}_{m}$, and
$C^{(m,k,\ell)}=|\det\left(J_{\mathcal{F}_{m}}\right)|J_{\mathcal{F}_{m}}^{-1}\left[\mu\left(\delta_{k,\ell}I_{3}+\underline{e}_{\ell}\underline{e}_{k}^{T}\right)+\lambda\underline{e}_{k}\underline{e}_{\ell}^{T}\right]J_{\mathcal{F}_{m}}^{-T}.$ (3.13)
@@
Here $\delta_{k,\ell}$ is the Kronecker delta. Note that $\widehat{v}_{ptc}^{(m)},\widehat{w}_{ptc}^{(m)}\in\widehat{\boldsymbol{\mathcal{S}}}_{ptc,0}^{(m)}$. Note that if $\text{supp}\left(v_{h}\right)\cap\text{supp}\left(w_{h}\right)\cap\Omega^{(m)}$ has an empty interior, at least one between $\widehat{v}_{ptc}^{(m)}$ and $\widehat{w}_{ptc}^{(m)}$ is the null function, and therefore $\boldsymbol{v}^{T}\mathbf{A}^{(i,j,k,\ell)}\boldsymbol{w}=0$.
We consider a low-rank approximation of the entries of $C^{(m,k,\ell)}$, i.e.
$C^{(m,k,\ell)}_{\alpha,\beta}\left(\xi_{1},\xi_{2},\xi_{3}\right)\approx\widetilde{C}^{(m,k,\ell)}_{\alpha,\beta}\left(\xi_{1},\xi_{2},\xi_{3}\right)=\sum_{r_{1},r_{2},r_{3}}\mathfrak{c}_{r_{1},r_{2},r_{3}}^{(m,k,\ell,\alpha,\beta)}c^{(m,k,\ell,\alpha,\beta)}_{(1,r_{1})}\left(\xi_{1}\right)c^{(m,k,\ell,\alpha,\beta)}_{(2,r_{2})}\left(\xi_{2}\right)c^{(m,k,\ell,\alpha,\beta)}_{(3,r_{3})}\left(\xi_{3}\right),$ (3.14)
for $\alpha,\beta=1,2,3,$ and we define $\widetilde{\mathbf{A}}^{(i,j,k,\ell)}$ by replacing $C^{(m,k,\ell)}$ with $\widetilde{C}^{(m,k,\ell)}$ into (3.12), i.e.
$\boldsymbol{v}^{T}\widetilde{\mathbf{A}}^{(i,j,k,\ell)}\boldsymbol{w}=\int_{[0,1]^{3}}\left(\nabla\widehat{v}_{ptc}^{(m)}\right)^{T}\widetilde{C}^{(m,k,\ell)}\;\nabla\widehat{w}_{ptc}^{(m)}\;\mathrm{d}{\underline{\xi}}.$ (3.15)
@@
Thanks to (3.14) and to the tensor structure of the spline space $\widehat{\boldsymbol{\mathcal{S}}}_{ptc,0}^{(m)}$, we have
$\widetilde{\mathbf{A}}^{(i,j,k,\ell)}=\sum_{r_{1},r_{2},r_{3}}\sum_{\alpha,\beta=1}^{3}\mathfrak{c}_{r_{1},r_{2},r_{3}}^{(m,k,\ell,\alpha,\beta)}\widetilde{A}^{(i,j,k,\ell,\alpha,\beta)}_{(3,r_{3})}\otimes\widetilde{A}^{(i,j,k,\ell,\alpha,\beta)}_{(2,r_{2})}\otimes\widetilde{A}^{(i,j,k,\ell,\alpha,\beta)}_{(1,r_{1})},$ (3.16)
@@
There is of course more than one approach to compute the low-rank approximation (3.14). A suitable possibility, proposed in [32], is to interpolate the left-hand side function using a coarse spline space and then performing a High Order Singular Value Decomposition (HOSVD) [14]. In this work, however, we follow [34] and employ the strategy described in [17, 15], which is based on interpolation with Chebyshev polynomials. Note that approximation (3.14) has to be computed only for $\alpha\geq\beta$ and $k\geq\ell$, because of the symmetries in the coefficients.
@@
where $\widehat{f}_{k}^{(i)}=\det\left(J_{\mathcal{G}_{i}}\right)\left(f_{k}\circ\mathcal{G}_{i}\right)$ and $\widehat{v}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}\in\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}$. We consider a low-rank approximation
$\widehat{f}_{k}^{(i)}\left(\xi_{1},\xi_{2},\xi_{3}\right)\approx\widetilde{f}_{k}^{(i)}\left(\xi_{1},\xi_{2},\xi_{3}\right)=\sum_{r_{1},r_{2},r_{3}}\mathfrak{f}^{(i,k)}_{r_{1},r_{2},r_{3}}f^{(i,k)}_{(1,r_{1})}\left(\xi_{1}\right)f^{(i,k)}_{(2,r_{2})}\left(\xi_{2}\right)f^{(i,k)}_{(3,r_{3})}\left(\xi_{3}\right),$ (3.18)
We define $\widetilde{\boldsymbol{f}}^{(i,k)}$ by replacing $\widehat{f}_{k}^{(i)}$ with $\widetilde{f}_{k}^{(i)}$ into (3.17), i.e.
$\boldsymbol{v}^{T}\widetilde{\boldsymbol{f}}^{(i,k)}=\int_{[0,1]^{3}}\widetilde{f}_{k}^{(i)}\;\widehat{v}_{sub}^{(i)}(\underline{\xi})\;\mathrm{d}\underline{\xi}=\sum_{r_{1},r_{2},r_{3}}\mathfrak{f}^{(i,k)}_{r_{1},r_{2},r_{3}}\int_{[0,1]^{3}}f^{(i,k)}_{(1,r_{1})}\left(\xi_{1}\right)\;f^{(i,k)}_{(2,r_{2})}\left(\xi_{2}\right)\;f^{(i,k)}_{(3,r_{3})}\left(\xi_{3}\right)\;\widehat{v}_{sub}^{(i)}(\underline{\xi})\;\mathrm{d}\underline{\xi}.$
@@
We now discuss the technical details of the low-rank approximation (3.18) considered in this work.
Assuming that $\mathcal{G}_{i}$ is given by (3.11), we observe that
$\widehat{f}_{k}^{(i)}=\left\{\begin{array}[]{ll}2g^{(i_{i})}_{k}\left(\xi_{1},\xi_{2},2\xi_{3}\right)&\text{if }0\leq\xi_{3}\leq\frac{1}{2},\\ 2g^{(i_{2})}_{k}\left(\xi_{1},\xi_{2},2\xi_{3}-1\right)&\text{if }\frac{1}{2}<\xi_{3}\leq 1.\end{array}\right.$
where $g^{(m)}_{k}=\det\left(J_{\mathcal{F}_{m}}\right)\left(f_{k}\circ\mathcal{F}_{m}\right)$ for $m=i_{1},i_{2}$. Note in particular that $\widehat{f}_{k}^{(i)}$ could be discontinuous at $\xi_{3}=\frac{1}{2}$. We use the Chebfun3F algorithm [15] to independently compute low-rank approximations of $g^{(i_{1})}_{k}$ and $g^{(i_{2})}_{k}$, i.e.
$\widetilde{g}^{(m)}_{k}\approx\det\left(J_{\mathcal{F}_{m}}\right)\left(f_{k}\circ\mathcal{F}_{m}\right)$ (3.19)
for $m=i_{1},i_{2}$. We then consider
$\widetilde{f}_{k}^{(i)}=\left\{\begin{array}[]{ll}2\widetilde{g}^{(i_{i})}_{k}\left(\xi_{1},\xi_{2},2\xi_{3}\right)&\text{if }0\leq\xi_{3}\leq\frac{1}{2},\\ 2\widetilde{g}^{(i_{2})}_{k}\left(\xi_{1},\xi_{2},2\xi_{3}-1\right)&\text{if }\frac{1}{2}<\xi_{3}\leq 1.\end{array}\right.$ (3.20)
Note that $\widetilde{f}_{k}^{(i)}$ can be easily written as a low-rank function whose multilinear rank is the sum of the multilinear ranks of $g^{(i_{1})}_{k}$ and $g^{(i_{2})}_{k}$. Therefore, as final step we reduce the rank of $\widetilde{f}_{k}^{(i)}$ using a purely algebraic approach based on the HOSVD. This truncation is performed with a relative tolerance close to machine precision, in order to remove any redundant terms while at the same time guaranteeing that(3.20) still holds. In particular, as it will be useful in the following, we observe that (3.20) yields
$\boldsymbol{v}^{T}\widetilde{f}^{(i,k)}=\int_{[0,1]^{3}}\widetilde{f}_{k}^{(i)}(\underline{\xi})\;\widehat{v}_{sub}^{(i)}(\underline{\xi})\;\mathrm{d}\underline{\xi}=\int_{[0,1]^{3}}\widetilde{g}_{k}^{(i_{1})}(\underline{\xi})\;\widehat{v}_{ptc}^{(i_{1})}(\underline{\xi})\;+\;\widetilde{g}_{k}^{(i_{2})}(\underline{\xi})\;\widehat{v}_{ptc}^{(i_{2})}(\underline{\xi})\;\mathrm{d}\underline{\xi},$ (3.21)
where $\widehat{v}_{ptc}^{(m)}=v_{h}\circ\mathcal{F}_{m}$, for $m=i_{1},i_{2}$.
@@
$\begin{split}\boldsymbol{v}^{T}\mathbf{A}^{(i,i,k,\ell)}\boldsymbol{w}&=\int_{\Theta^{(i)}}2\mu\;\varepsilon(\underline{e}_{k}v_{h}):\varepsilon(\underline{e}_{\ell}w_{h})+\lambda\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\left(\nabla\cdot\underline{e}_{\ell}w_{h}\right)\mathrm{d}{\underline{x}}=\\ &=\int_{[0,1]^{3}}\left(\nabla\widehat{v}_{sub}^{(i)}\right)^{T}Q^{(i,k,\ell)}\nabla\widehat{w}_{sub}^{(i)}\mathrm{d}{\underline{\xi}},\end{split}$ (3.22)
where $\widehat{v}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}$, $\widehat{w}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}\in\widehat{\boldsymbol{\mathcal{S}}}^{(i)}_{sub,0}$, and